OpenStudy (anonymous):

derivative of cos(5-theta)

OpenStudy (surry99):

use chain rule

OpenStudy (anonymous):

so would my answer be sin(theta-5) ??

OpenStudy (surry99):

not quite...why did you reverse the (5 - theta)

OpenStudy (anonymous):

becuase... cos(5-theta) (d/dx)=-sin(5-theta)(-1) (d/dx)=sin(theta-5)

OpenStudy (anonymous):

That's valid since \(\sin x\) is an odd function; i.e. \(\sin (-x)=-\sin x\). What @gymkara did was reverse this property by noting that \(-\sin(5-\theta) = \sin(-(5-\theta)) = \sin(\theta-5)\).

OpenStudy (anonymous):

however, there's still a negative sign missing in his derivative!

OpenStudy (surry99):

exactly!

OpenStudy (anonymous):

I just realized that... >_>

OpenStudy (anonymous):

so should it be -sin(theta-5) ?

OpenStudy (anonymous):

You didn't apply chain rule! \((\cos(5-\theta))^{\prime} = -\sin(5-\theta)\cdot{\color{red}{(5-\theta)^{\prime}}} =\ldots\)

OpenStudy (surry99):

I don't use identities unless I need to so my answer would have been: -sin(5-theta)*-1 = sin(5 - theta)

OpenStudy (anonymous):

@ChristopherToni if I take the derivative of the part in red, it's -1...so if I multiply -sin(5-theta) by -1, I get sin(theta-5)

OpenStudy (anonymous):

\(-\sin(5-\theta) \cdot (-1) = \sin(5-\theta)\), not \(-\sin(5-\theta)\).

OpenStudy (anonymous):

Oh DUH!!! I'm stupid!! Thank you :)

OpenStudy (anonymous):

No worries; at least you see where you were going wrong - that's the important part! :-)